OpenStudy (anonymous):

x-6y-2z=-8 -x+5y+3z=2 3x -2y -4z=18 can you solve this using substitution method please i have my finale tom and i need to see how its done

OpenStudy (anonymous):

please help i have like ten problems on my final for this and i totally forgot

OpenStudy (radar):

Decide which variable you wish to substitute, x, y or z ?

OpenStudy (anonymous):

i know the answers i just need to know how to do them

OpenStudy (anonymous):

http://www.sosmath.com/soe/SE3/SE3.html thats a tutorial about it.

OpenStudy (radar):

I was going to walk you thru but I wanted you to be the decision maker on which variable you want to substitute.

OpenStudy (anonymous):

um it doesnt matter x

OpenStudy (radar):

I would suggest x, but that is just me, and you are you.

OpenStudy (anonymous):

i picked x

OpenStudy (radar):

Good choice. in the first equation x-6y-2z=-8. we need to isolate the x, in an expression that gives x in terms of y and z. Follow this sequence: x-6y-2z=-8 x=6y + 2z - 8 That has isolated x and shows x =6y+2z-8 Now substitute this expression for x in the two lower equations.

OpenStudy (anonymous):

how do you do the rest of it?

OpenStudy (radar):

-6y-2z+8+5y+3z=2 -y+z=-6 This is result of substitution in the second equation, you do it in the thrid equation, while I eat my pizza.

OpenStudy (radar):

Now substitute it in the third equation: 3(6y+2z-8)-2y-4z=18 18y+6z-24 -2y-4z=18 16y +2z=18+24 16y+2z=42

OpenStudy (radar):

Here is what we have now: x=6y+2z-8 -y+z=-6 16y+2z=42 Note the last two equation have only two unknowns (y and z) Take the -y+z=-6 and isolate the z getting z=y-6 and substitute this for z in the equation 16y + 2z=42. You would get 16y+2(y-6)=42 16y+2y-12=42 18y=54 y=3 Yeah! one down now substitute 3 for y in 16y + 2z = 42 getting 48 +2z =42 2z=-6 z=-3 (another one down), substitute for both y and z in x=6y+2z-8 x= 18-6-8=4 x=4

OpenStudy (radar):

I kind of wanted to verify this: x=4, z=-3, and y=3 x-6y-2z=-8 4-18+6=-8 -18+10=-8 -8=-8 Good enuff for me.